OpenStudy (anonymous):

lim as x approaches negative infinity (x + (x^2 +x+1)^0.5)

OpenStudy (solomonzelman):

\(\large\color{black}{ }\)\[\LARGE \lim_{x \rightarrow \infty }~x+\sqrt{x^2+x+1}\]Right?

OpenStudy (solomonzelman):

maybe it is\[\LARGE \lim_{x \rightarrow \infty }~x\color{red}{-}\sqrt{x^2+x+1}~~~?\]

OpenStudy (anonymous):

its plus

OpenStudy (anonymous):

to negative infinity

OpenStudy (solomonzelman):

Ohh, my bad

OpenStudy (solomonzelman):

\[\LARGE \lim_{x \rightarrow ~-\infty }~x+\sqrt{x^2+x+1}\]

OpenStudy (solomonzelman):

Sorry about that.

OpenStudy (solomonzelman):

\[\LARGE \lim_{x \rightarrow ~-\infty }~x+\sqrt{x^2+x+1} \] \[\Large \lim_{x \rightarrow ~-\infty }~(x+\sqrt{x^2+x+1} )\frac{-x+\sqrt{x^2+x+1}}{-x+\sqrt{x^2+x+1}}\]

OpenStudy (solomonzelman):

Rationalizing...

OpenStudy (solomonzelman):

\[\Large \lim_{x \rightarrow ~-\infty }~\frac{x+1}{-x+\sqrt{x^2+x+1}} \]

OpenStudy (solomonzelman):

Then re-write the sum of the 2 limits (hope you know which) separately.

OpenStudy (anonymous):

re-write by dividing by x ?

OpenStudy (solomonzelman):

I meant \[\large \lim_{x \rightarrow ~-\infty }~\frac{x}{-x+\sqrt{x^2+x+1}} +\large \lim_{x \rightarrow ~-\infty }~\frac{1}{-x+\sqrt{x^2+x+1}} \]

hartnn (hartnn):

you could be done if you divide numerator and denominator by x in this step itself \(\Large \lim_{x \rightarrow ~-\infty }~\frac{x+1}{-x+\sqrt{x^2+x+1}}\)

OpenStudy (solomonzelman):

yes, likely...

OpenStudy (solomonzelman):

yes, likely...

OpenStudy (anonymous):

so 1 / (-1 -inifinity) + zero = zero

OpenStudy (solomonzelman):

no, I don't think this is right,

hartnn (hartnn):

\(\Large \lim_{x \rightarrow ~-\infty }~\frac{x+1}{-x+\sqrt{x^2+x+1}} \\= \Large\Large \lim_{x \rightarrow ~-\infty }~\frac{1+1/x}{-1+\sqrt{1+1/x+1/x^2}} \) since x -> - infinity 1/x ->0 1/x^2 ->0 so you can simply plug 0 in place of 1/x and /x^2

OpenStudy (anonymous):

ah 1/(1+1) = 1/2

OpenStudy (solomonzelman):

No it is not 1/2 I was about to reply, but as again latex disconnected me.

OpenStudy (solomonzelman):

I lost all of it:)

OpenStudy (solomonzelman):